which value is the largest? 11/5, 2.05, 2.5, 9/4
Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

11/5=2.2
9/4=2.25
2.5 is largest

Answer by mananth(16949)

11/5 = 2.2,
9/5= 2.25

2.5 is the largest